Walt made an extra$10000 last year from a part time job. He invested part of the money at 10% and the rest at 8%. He made a total of $940 in interest. How much was invested at 8%?

Answer by Boreal(15235)

x=8% return
10000-x=10% return
.08x+ .1 (10,000-x)=940
.08x + 1000-.1x=940 ;; collect x terms and subtract 1000 from both sides
-0.02x=-60 ;; divide both sides by -0.02
x=3000
3000 at 8%=$240
7000 at 10%=$700
$940 total